OpenStudy (anonymous):

how is 80/4*2-2*2's answer 36 please show steps

OpenStudy (saifoo.khan):

Remember oder of operations. BoDMAS Brackets Divide Multiply Add Subtract.

OpenStudy (saifoo.khan):

Now try it out.

OpenStudy (anonymous):

80/4 = 20 20*2 = 40 2*2 = 4 40 - 4 = 36 :)